Henry can build 2 birdhouses in 30 minutes. How many birdhouses can he build in four hours?

Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

Henry can build 2 birdhouses in 30 minutes. How many birdhouses can he build in four hours?
-------
Use a proportion::
houses/time = houses/time
------
h/4 hrs = 2/(1/2 hrs)
------
# of houses = 8/(1/2) = 16
-------
Cheers,
Stan H.
---------

Answer by jim_thompson5910(35256)

30 minutes = 0.5 hours


The idea is to form a proportion and then isolate the variable.


(2 birdhouses)/(0.5 hours) = (x birdhouses)/(4 hours)


2/0.5 = x/4


2*4 = 0.5*x ... cross multiply


8 = 0.5x


0.5x = 8


0.5x/0.5 = 8/0.5 ... divide both sides by 0.5


x = 16


He can make 16 birdhouses.